OpenStudy (anonymous):

How come when you square root a number, like for example 9, its + and - 3

OpenStudy (anonymous):

let x=sqrt(9) or xsquaring both sides we have x^2=9 x^2-9=0 x^2 -(3)^2=0 (x+3)(x-3)=0 (factorising) thus either x+3=0(means x=-3) or x-3=0(meansx=3) hence x=+3 or -3

OpenStudy (anonymous):

Thank you :) I get it now!

OpenStudy (anonymous):

WELCOME

OpenStudy (accessdenied):

A small note about square roots: The symbol \(\sqrt{~~}\) (the radical sign) is only defined as the principal/positive square root. All nonnegative real numbers have two square roots: \(\pm \sqrt{N} \to \sqrt{N} \text{ and } \neg \sqrt{N} \). Sometimes there is confusion in the symbol and the actual term "square root."